OpenStudy (anonymous):

Factor completely: 3bx2 − 9x3 − b + 3x

OpenStudy (anonymous):

Right, I take it for -9x^3+3bx^2+3x-b=0 I guess?

OpenStudy (anonymous):

So, here's what I do. -3x(3x^2-1) + b( 3x^2-1) = 0 => (b-3x)(3x^2-1)=0

OpenStudy (anonymous):

By now, the answer should be obvious.

Directrix (directrix):

@NeetziD If we are factoring, then why set the expression equal to zero as if we were finding the roots? Would the solution not end here: --> (b-3x)(3x^2-1) >>Factor completely: 3bx2 − 9x3 − b + 3x

OpenStudy (anonymous):

@Directrix Sorry... took it for a equation. =P didn't notice the factor bit... dumbed down there

Directrix (directrix):

Hey, no need to be sorry about anything. I was just asking. And, on that open cone question the Asker said is wrong, I got the same thing you did.